OpenStudy (anonymous):

Solve by using elimination 2x-5y= 11 4x+10y= 18

OpenStudy (jdoe0001):

hmmm lemme redo that

OpenStudy (jdoe0001):

\(\large \begin{array}{llcll} 2x-5y= 11&{\color{brown}{ \times -2}}\implies &\cancel{-4xx }+10y=-22\\ 4x+10y= 18&&\cancel{ 4x }+10y=18 \\\hline\\ &&\square ?+\square ?=\square ? \end{array} \)

OpenStudy (jdoe0001):

so...what does that give you?

OpenStudy (anonymous):

Isn't that 0= -40 ?

OpenStudy (jdoe0001):

0? whatever happened to the "y"?

OpenStudy (anonymous):

Aren't you supposed to subtract the bottom equation from the top one?